# Neural Spectro-polarimetric Fields

### v1

@@ 4. Neural Spectro-Polarimetric Field
We define the spectro-polarimetric field $f(\cdot)$, which describes the Stokes vector of a light ray at a given position, direction, and wavelength:$\mathbf{s}=f(x,y,z,\theta,\phi,\lambda)$, where $\mathbf{s}$ is the Stokes vector. Spectro-polarimetric field can be thought of as an extended parameterization of plenoptic function to account for polarization. Unlike the simplified plenoptic function used in NeRF [Mildenhall et al., 2020] shown in Figure 3(a), the spectro-polarimetric field incorporates wavelength as an input variable and the Stokes vector as an output, as depicted in Figure 3(b).
@@
Using NeSpoF, we can render the Stokes vector $\mathbf{s}_{\text{rendered}}$ from a ray $\mathbf{r}$ emitted from a particular pixel at a wavelength $\lambda$ by applying volumetric rendering as follows:
@@
4.1. Modeling Physically-valid Spectral Stokes Vector
We design NeSpoF with a positional MLP, $F_{\Theta}^{p}$, and a spectro-directional MLP, $F_{\Theta}^{d}$, as shown in Figure 3(c). Although the positional MLP is based on NeRF, we provide a brief description for clarity. The positional MLP estimates the volumetric density $\sigma$ at the position $(x,y,z)$ and extracts the positional Stokes-vector feature $f^{p}$. The positional encoding $\gamma_{k}(x)$ [Tancik et al., 2020] is employed to learn high-frequency details of the Stokes-vector distribution with respect to the position:
@@ 4.1. Modeling Physically-valid Spectral Stokes Vector
Wavelength as input
We use wavelength $\lambda$ as an input to the spectro-directional MLP in addition to the direction variables $\theta$ and $\phi$, unlike NeRF that takes only the direction variables as input and outputs as three RGB radiance values. In this way, we can query the Stokes vector from NeSpoF at an arbitrary wavelength $\lambda$, suitable for spectral analysis and measurements of the Stokes vectors. We apply the positional encoding of Equation (4) to the wavelength with the hyperparameter value $k=1$. We can control $k$ for representing higher frequency changes of the spectral Stokes vector.
@@
It is crucial to ensure that NeSpoF outputs physically-valid Stokes vectors. As described in Section 3, a Stokes vector is physically valid if the following condition is met: $s_{0}^{2}\geq s_{1}^{2}+s_{2}^{2}+s_{3}^{2}$. As a naïve approach of modeling the output of the spectro-directional MLP as Stokes-vector elements fails to guarantee the physical validity of the output Stokes vector, we propose to use intermediate polarimetric properties as outputs: $X_{0}$, $X_{1}$, $X_{2}$, $X_{3}$. We then map the values to the total intensity $s_{0}$, DoP $\rho$, ellipticity $\chi$, and azimuth angle $\psi$ as follows:
@@
Figure 4. (a) We built an experimental prototype with a monochromatic camera, LCTF, and QWP mounted on a rotation stage. (b) We capture images with the rotating QWP by varying the angle $\theta$ at each target wavelength. (c) To calibrate the spatially-varying spectro-polarimetric variation of the prototype, we use 16 samples with known Stokes vectors made by the polarization-state generator. (d) The captured images with varying QWP angles clearly exhibit the spatially-varying spectro-polarimetric characteristics. (e) After calibration, we achieve accurate polarimetric reconstruction of the linearly-polarized at $0^{\circ}$ degree.
5. Multi-view Hyperspectral-polarimetric Imaging and Dataset
To test NeSpoF on real-world scenes, we construct a portable imaging system. We propose a calibration method to address the spatially-varying imperfections of the experimental prototype. Using the imaging system, we capture and release the first real-world multi-view hyperspectral-polarimetric image dataset.
5.1. Hyperspectral-polarimetric Imaging
@@
Our experimental prototype, shown in Figure 4, consists of a monochromatic sensor equipped with a 35 mm lens, a liquid crystal tunable filter (LCTF), and a quarter-wave plate (QWP) mounted on a motorized rotation stage. The LCTF can adjust spectral transmission. We placed a cut-off filter for the IR and UV wavelengths in front of the LCTF.
@@
Even though we designed our imaging system using professional optical devices, the experimental prototype still presents deviation from a commonly-assumed model for the polarization modulation of LCTF, which has often been assumed to be an ideal linear polarizer [Fan et al., 2023]. To overcome this problem, we propose to optimize for spatially-varying matrices $\mathbf{M}(\lambda,p;\lambda)$ that act as polarization modulation by the LCTF, where $p$ is a pixel. To this end, we capture 16 samples with known Stokes vectors constructed with an LP, a QWP, and a collimated LED light source. We then solve for the matrix $\mathbf{M}$ as follows:
@@
where $I^{\lambda,\theta_{k}}_{\mathrm{meas},i}(p)$ is the intensity measurement of the $i$-th sample for the pixel $p$ at the rotation angle $\theta_{k}$ of the QWP. $\mathbf{s}_{\text{u}}$ is the Stokes vector of unpolarized light. We used four uniformly-sampled $\theta$ in the $[0,\pi]$ range for this calibration. We refer to the Supplemental Document for its normal-equation formulation.
@@
Once the matrix $\mathbf{M}$ is obtained, we can reconstruct the Stokes vector $\mathbf{s}^{\lambda}_{\text{meas}}$ for a pixel $p$ of an arbitrary scene by solving the least-squares problem:
(12) $\mathbf{s}_{\text{meas}}(\lambda,p)=\mathop{\text{argmin}}\limits_{\mathbf{s}^{\lambda}}\sum\limits_{k=1}^{K}{\left(I^{\lambda,\theta_{k}}_{\mathrm{meas}}(p)-c(\lambda)\mathbf{M}(\lambda,p;\lambda){\mathbf{Q}}(\theta_{k},\lambda)\mathbf{s}(\lambda,p)\right)}^{2},$
where $c(\lambda)=T(\lambda)F(\lambda)$. We refer to the Supplemental Document for its normal-equation formulation. Figure 4(c-e) demonstrates the calibration setup and reconstruction of the Stokes vector of ${}^{\circ}$ linearly-polarized light.
5.2. Multi-view Hyperspectral-polarimetric Dataset
@@
Using the experimental prototype, we acquired the first multi-view hyperspectral-polarimetric dataset of four real-world scenes. At each camera viewpoint, we capture raw measurements and reconstruct the per-pixel Stokes vector $\mathbf{s}^{\lambda}_{\text{meas}}$ using Equation (12). It takes roughly 10 seconds to capture a hyperspectral-polarimetric image with 21 spectral channels at four QWP angles. We then perform local-to-world coordinate conversion to use it for NeSpoF. We obtain the geometric parameters of the multiple viewpoints and the camera intrinsic parameters using a structure-from-motion method, COLMAP [Schönberger and Frahm, 2016], for which we use the RGB intensity images obtained by applying the spectrum-to-RGB conversion as well as the gamma correction. We refer to the Supplemental Document for more details about the dataset, calibration, and coordinate conversion.
Synthetic Scenes
We render multi-view hyperspectral-polarimetric images of four synthetic scenes using Mitsuba3 [Jakob et al., 2022]. We manually assign the multispectral polarimetric BRDFs [Baek et al., 2020] to meshes [Mildenhall et al., 2020; Bitterli, 2016]. Since only five spectral channels are available (450, 500, 550, 600, and 650 nm) in the BRDF dataset, we fit a fourth-order polynomial function at each rendered pixel along the spectrum.
6. Assessment
We train NeSpoF using multi-view hyperspectral polarimetric images for each scene with a batch of rays with the size 32$\times$32$\times$4 and 200k training iterations, which takes 10 hours on an NVIDIA GeForce RTX 3090 GPU. We use a learning rate of $5\times 10^{-4}$.
6.1. View-spectrum-polarization Synthesis
Qualitative Evaluation on Synthetic Scenes
Figure 6(a) shows the rendered hyperspectral-polarimetric Stokes-vector images at two novel views of a synthetic scene. The synthesized results accurately match the ground-truth images as visualized in AoLP, DoP, and ToP. In particular, NeSpoF is capable of synthesizing unique spectro-polarimetric characteristics of the metallic dragon, spheres, and teapots made of different materials. Please see the Supplemental Document for additional synthetic-scene results.
Quantitative Evaluation on Synthetic Scenes
Figure 6(b) shows the quantitative results of NeSpoF on the test views of four synthetic scenes. In terms of wavelength, we observe a PSNR of 35 dB and an RMSE of 0.020 at a wavelength of 650 nm, with better reconstructions for other wavelengths. For each Stokes-vector element, the radiance component, which usually exhibits the most complex variation, has a PSNR of 33 dB and RMSE of 0.023, while the other elements demonstrate higher PSNRs and lower RMSEs. The scene-dependent quantitative analysis shows a minimum PSNR of 30.4 dB.
Figure 5. (a) NeSpoF enables rendering of hyperspectral Stokes vectors from novel views and (b) facilitates the analysis of the hyperspectral intensity of unpolarized diffuse and polarized components, which include specular reflections and polarized inter-reflections. In the synthetic dragon scene, we analyze diffuse and specular intensity at each wavelength and observe polarized inter-reflections between walls. For the real-world scene, we examine the hyperspectral intensity of a flower and its reflection on glass, both showing different polarization states depending on reflection angles.
Real-world Scenes
@@ 6.1. View-spectrum-polarization Synthesis
Figure 10 presents the rendered images of NeSpoF on two real-world scenes. The first scene encompasses a laptop display and a mobile phone equipped with a protective film. In front of the mobile phone is a linear polarizer. While the laptop display emits linearly polarized light, the light from the mobile phone display interacts with the stress-affected protective film, leading to complex spectro-polarimetric patterns. The second scene features a CD demonstrating unique spectro-polarimetric characteristics attributed to diffraction caused by micro-scale surface structures, paired with a doll covered by a linear-polarizer film. As shown in the AoLP and ToP visualization at two distinct wavelengths, NeSpoF accurately synthesizes the invisible information, matching with the ground-truth data. Moreover, NeSpoF effectively regularizes low SNR measurements from hyperspectral-polarimetric images, resulting in high-fidelity synthesis of hyperspectral Stokes fields in real-world scenes. Due to the low SNR hyperspectral-polarimetric measurements, quantitative evaluation has been excluded. Figure 8 shows the rendered images with NeSpoF, exhibiting view-dependent hyperspectral-polarimetric information oh the reflections on crystal and a baseball inside of a glass box. Figure 7 demonstrates novel view synthesis by NeSpoF for an outdoor scene under sky lighting. This allows for analysis of the spectro-polarimetric properties of sky light, which is partially linearly polarized, as indicated by high values of $s_{1}$ and $s_{2}$. The illumination spectrum of the sky can also be extracted, as depicted in Figure 7(b). This enables rendering of the hyperspectral intensity at a novel view and relighting by replacing the original illumination with the target illumination, here CIE D65 illuminant, shown in Figure 7(c).
@@
Figure 5 shows the analysis of the diffuse and specular hyperspectral intensity with NeSpoF. Here, we modeled diffuse reflection as unpolarized intensity, while specular reflection is modeled as polarized intensity. This enables us to examine the hyperspectral intensity of each diffuse and specular component separately, at each wavelength. Within these separated hyperspectral diffuse and specular components, we can observe more complex light transport phenomena such as the polarized inter-reflection between walls for the dragon scene. For the real-world scene, it is observed that the diffuse component of the flower is unpolarized, and hence not visible in the specular image. However, its reflection on the glass is clearly visible in the specular image due to the strong polarization caused by Fresnel reflection near a Brewster angle. When rendering the frontal-view specular image of the glass, the strong specular signal of the flower on the glass is not visible, as the imaging geometry deviates from the Brewster-angle configuration.
6.2. Ablation Study
Being the first neural model to address the modeling of hyperspectral Stokes fields within a scene, we here focus on the analysis of NeSpoF. Comparative studies can be found in the Supplemental Document.
@@ 6.2. Ablation Study
We test the weight of the loss function $w_{i}$ for the i-th Stokes vector in Equation (8). Using the mean-squared-error (MSE) loss with the uniform weights of $w_{i}=1$ results in unstable training for the Stokes-vector elements with low values, e.g. the last element $[\mathbf{s}]_{3}$. Figure 9 shows that using our weighted loss of Equation (8) allows for accurate reconstruction of the original Stokes vector by balancing the scale difference between the Stokes-vector elements.
@@
We test the impact of positional-encoding frequency $k$ of Equation (4) for the input wavelength $\lambda$. Figure 9 shows that $k=1$ provides a high-fidelity plenoptic synthesis as visualized in the AoLP images.

### v2

@@ 4. Neural Spectro-polarimetric Field
We define spectro-polarimetric field $f(\cdot)$ as a description of the Stokes-vector field at any given position, direction, and wavelength: $\mathbf{s}=f(x,y,z,\theta,\phi,\lambda)$, where $\mathbf{s}$ is the Stokes vector. Unlike the simplified plenoptic function used in NeRF [Mildenhall et al., 2020], the spectro-polarimetric field incorporates wavelength as an input variable and outputs the Stokes vector. See Figure 3.
@@
Using NeSpoF, we can render the Stokes vector $\mathbf{s}_{\text{rendered}}$ from a ray $\mathbf{r}$ emitted from a particular camera pixel at a wavelength $\lambda$ by applying volumetric rendering as follows:
@@
Model Design
We design NeSpoF with a positional MLP, $F_{\Theta}^{p}$, and a spectro-directional MLP, $F_{\Theta}^{d}$, as shown in Figure 3(c). The positional MLP adopts the original NeRF design. It estimates the volumetric density $\sigma$ at the position $(x,y,z)$ and extracts the positional Stokes-vector feature $f^{p}$. The positional encoding $\gamma_{k}(x)$ [Tancik et al., 2020] enables learning high-frequency details of the Stokes-vector distribution with respect to the position:
@@
Wavelength as an Input
We use wavelength $\lambda$ as an input to the spectro-directional MLP in addition to the direction variables $\theta$ and $\phi$, unlike NeRF that takes only the direction variables as input and outputs as three RGB radiance values. In this way, we can query the Stokes vector from NeSpoF at an arbitrary wavelength $\lambda$, suitable for spectral analysis and measurements of the Stokes vectors. We apply the positional encoding of Equation (4) to the wavelength with the hyperparameter value $k=1$.
@@
It is crucial to ensure that NeSpoF outputs physically-valid Stokes vectors. A Stokes vector is physically valid if the following condition is met: $s_{0}^{2}\geq s_{1}^{2}+s_{2}^{2}+s_{3}^{2}$ [Collett, 2005]. A naïve approach of modeling the output of the spectro-directional MLP as Stokes-vector elements fails to guarantee the physical validity of the output Stokes vector. Hence, we propose to use intermediate polarimetric properties as outputs: $X_{0}$, $X_{1}$, $X_{2}$, $X_{3}$. We then map the values to the total intensity $s_{0}$, DoP $\rho$, ellipticity $\chi$, and azimuth angle $\psi$ as follows:
@@
5. Spectro-polarimetric Imaging
To test NeSpoF on real-world scenes, we construct an imaging system and propose a calibration method to address the imperfections of the imaging system. Using the imaging system, we capture and release the first real-world multi-view hyperspectral-polarimetric image dataset.
Figure 4. (a) Experimental prototype and (b) schematic diagram. We capture images with the QWP at four different angles at each wavelength. (c) The spatially-varying spectro-polarimetric transmission of the LCTF results in spatial non-uniformity. The images of linearly-polarized uniform light are shown at the four QWP angles. (d) We calibrate such non-uniformity with spatially-varying Mueller matrices, resulting in accurate polarimetric reconstruction of $0^{\circ}$ degree linearly-polarized light.
5.1. Imaging
@@
Our experimental prototype, shown in Figure 4(a), consists of a monochromatic sensor equipped with a 35 mm lens, a liquid crystal tunable filter (LCTF), and a quarter-wave plate (QWP) mounted on a motorized rotation stage. The LCTF can adjust spectral transmission. We placed a cut-off filter for the IR and UV wavelengths in front of the LCTF.
@@
Figure 5. Synthetic evaluation. NeSpoF achieves an accurate synthesis of hyperspectral Stokes vectors at novel views, evidenced in both qualitative and quantitative analysis.
Figure 6. Multi-view hyperspectral polarimetric dataset.
@@
Even though we designed our imaging system using professional optical devices, the experimental prototype still presents deviation from a commonly-assumed model for the polarization modulation of LCTF, which has often been assumed to be an ideal linear polarizer [Fan et al., 2023]. As shown in Figure 4(c), the captured intensity through our imaging system shows spatially non-uniform spectro-polarimetric variations. To overcome this problem, we propose to optimize for spatially-varying matrices $\mathbf{M}(\lambda,p;\lambda)$ that act as polarization modulation by the LCTF, where $p$ is a pixel. We capture 16 samples with known Stokes vectors constructed with a LP, a QWP, and a collimated LED light source. We then solve for the matrix $\mathbf{M}$ as follows:
@@
where $I^{\lambda,\theta_{k}}_{\mathrm{meas},i}(p)$ is the intensity measurement of the $i$-th sample for the pixel $p$ at the rotation angle $\theta_{k}$ of the QWP. $\mathbf{s}_{\text{u}}$ is the Stokes vector of unpolarized light. We used four uniformly-sampled $\theta$ in the $[0,\pi)$ range for this calibration. We refer to the Supplemental Document for its normal-equation formulation.
@@
Once the matrix $\mathbf{M}$ is obtained, we can reconstruct the per-pixel Stokes vector $\mathbf{s}_{\text{meas}}$ for a wavelength $\lambda$ by solving the least-squares problem:
(12) $\mathop{\text{minimize}}\limits_{\mathbf{s}_{\text{meas}}(\lambda)}\sum\limits_{k=1}^{K}{\left(I^{\lambda,\theta_{k}}_{\mathrm{meas}}-c(\lambda)\mathbf{M}(\lambda;\lambda){\mathbf{Q}}(\theta_{k},\lambda)\mathbf{s}_{\text{meas}}(\lambda)\right)}^{2},$
where $c(\lambda)=T(\lambda)F(\lambda)$. We exclude the pixel dependency in the equation for readability. Figure 4(d) shows the results of the calibrated Stokes-vector reconstruction.
Figure 7. We evaluate the impacts of (a) the utilization of weighted loss and (b) the frequency $k$ of the positional encoding for wavelength. Using the weighted loss and the frequency of $k=1$ enables better polarimetric reconstruction. We report PSNRs of the estimated ToP and CoP for the former, and DoP and AoLP for the later.
Figure 8. NeSpoF allows for synthesizing view-dependent hyperspectral-polarimetric patterns of the crystal and a baseball.
5.2. Dataset
Using the experimental prototype, we acquired the first multi-view hyperspectral polarimetric dataset of four real-world scenes. At each camera viewpoint, we capture raw measurements and reconstruct the per-pixel Stokes vector $\mathbf{s}^{\lambda}_{\text{meas}}$ using Equation (12). It takes roughly 10 seconds to capture a hyperspectral-polarimetric image with 21 spectral channels at four QWP angles ($-90^{\circ}$, $-45^{\circ}$, $30^{\circ}$, and $60^{\circ}$[Ambirajan and Look Jr, 1995]. We then perform local-to-world coordinate conversion to use it for NeSpoF. We obtain the geometric parameters of the multiple viewpoints and the camera intrinsic parameters using a structure-from-motion method, COLMAP [Schönberger and Frahm, 2016], for which we use the RGB intensity images obtained by applying the spectrum-to-RGB conversion as well as the gamma correction. To test NeSpoF on clean images without measurement noise, we also render multi-view hyperspectral polarimetric images of four synthetic scenes using Mitsuba3 [Jakob et al., 2022]. We manually assign the multispectral polarimetric BRDFs [Baek et al., 2020] to meshes [Mildenhall et al., 2020; Bitterli, 2016]. Since only five spectral channels are available (450, 500, 550, 600, and 650 nm) in the BRDF dataset, we fit a fourth-order polynomial function at each rendered pixel along the spectrum. Figure 6 shows the thumbnails of our multi-view hyperspectral polarimetric dataset. We refer to the Supplemental Document for more details about the dataset, calibration, and coordinate conversion.
6. Assessment
6.1. View-spectrum-polarimetric Synthesis
We train NeSpoF for each scene in the dataset and test the capability of NeSpoF on view-spectrum-polarization synthesis.
Synthetic Scenes
We demonstrate the effectiveness of NeSpoF on the synthetic dataset. Figure 5(a) shows the rendered and ground-truth hyperspectral-polarimetric Stokes-vector images. The synthesized results accurately match the ground truth such as visualized in AoLP, DoP, and ToP. Figure 5(b) shows the quantitative results on all test views of the synthetic dataset. In terms of wavelength, NeSpoF obtains PSNR values of 35 dB and an RMSE of 0.020 at 630 nm, with better reconstructions for other wavelengths. For each Stokes-vector element, the radiance component, which usually exhibits the most complex variation, has a PSNR of 33 dB and RMSE of 0.023, while the other elements demonstrate higher PSNRs and lower RMSEs. The scene-dependent quantitative analysis shows a minimum PSNR of 30.4 dB. Refer to the Supplemental Document for additional synthetic-scene results.
@@ 6.1. View-spectrum-polarimetric Synthesis
NeSpoF enables modeling complex view-spectro-polarimetric information of real-world scenes. Figure 8 demonstrates NeSpoF on synthesizing the intricate variations of AoLP from the reflections of the crystal and the baseball inside of a glass box. Figure 9 shows two real-world scenes. The first scene consists of a laptop and a mobile phone equipped with a protective film. In front of the mobile phone is a linear polarizer. The laptop display emits clean linearly polarized light whereas the light from the phone display interacts with the stress-affected protective film. The second scene contains a CD demonstrating unique spectro-polarimetric characteristics attributed to diffraction caused by micro-scale surface structures, paired with a doll covered by a linear polarizer. NeSpoF accurately models the resulting complex spectro-polarimetric patterns. Moreover, the regularization capability of NeSpoF can be observed in low SNR regions.
Figure 9. We demonstrate the synthesis of view-spectrum-polarization using NeSpoF in two real-world scenes that exhibit unique view-spectro-polarimetric characteristics, specifically a mobile display covered with stressed protection film, a linear polarizer, and a CD. We visualize AoLP at two different wavelengths from a novel view, and ToP at different views.
Figure 10. (a) NeSpoF facilitates the rendering of hyperspectral-polarimetric images of an outdoor scene at a novel view, revealing the partially-linear polarization state of skylight. (b) We extract the hyperspectral intensity of the skylight and (c) use it to relight the scene from two novel views under D65 illumination.
Novel-view Spectral Relighting
Figure 10 demonstrates novel view synthesis by NeSpoF for an outdoor scene under sky lighting. This allows for analysis of the spectro-polarimetric properties of sky light, which is partially linearly polarized, as indicated by high values of $s_{1}$ and $s_{2}$. The illumination spectrum of the sky can also be extracted, as depicted in Figure 10(b). This enables rendering of the hyperspectral intensity at a novel view and relighting by replacing the original illumination with the target illumination, here CIE D65 illuminant, shown in Figure 10(c).
@@
Figure 11 shows the analysis of the unpolarized and polarized hyperspectral intensity with NeSpoF. This enables us to separate not only diffuse and specular components but also more complex light transport, such as the polarized inter-reflection between walls for the synthetic dragon scene. For the real-world scene, it is observed that the diffuse component of the flower is unpolarized, and hence not visible in the specular image. However, its reflection on the glass is clearly visible in the specular image due to the strong polarization caused by Fresnel reflection near a Brewster angle. When rendering the frontal-view specular image of the glass, the strong specular signal of the flower on the glass is not visible, as the imaging geometry deviates from the Brewster-angle configuration.
6.2. Analysis of NeSpoF
Being the first neural model to address the modeling of hyperspectral Stokes fields, we here focus on the analysis of NeSpoF. Comparative studies can be found in the Supplemental Document.
@@ 6.2. Analysis of NeSpoF
We test the weight of the loss function $w_{i}$ for the i-th Stokes vector in Equation (8). Using the mean-squared-error (MSE) loss with the uniform weights of $w_{i}=1$ results in unstable training for the Stokes-vector elements with low values, e.g. the last element $[\mathbf{s}]_{3}$. Figure 7(a) shows that using our weighted loss of Equation (8) allows for accurate reconstruction of the original Stokes vector by balancing the scale difference between the Stokes-vector elements.
Figure 11. (a) NeSpoF enables the separation of unpolarized diffuse and polarized components, which include specular reflections and polarized inter-reflections for the synthetic scene. (b) For the real-world scene, we identify the different polarization states of a flower and its reflection to glass depending on reflection angles.
Figure 12. Input images with severe noise and fewer input views lead to performance degradation of NeSpoF.
@@
We test the impact of positional-encoding frequency $k$ of Equation (4) for the input wavelength $\lambda$. Figure 7(b) shows that $k=1$ provides a high-fidelity plenoptic synthesis as visualized in the DoP and AoLP images.
Measurement Noise
The performance of NeSpoF is affected by measurement noise, as shown in rendering artifacts of low-SNR regions in real-world scenes such as the black background in Figure 1. We evaluate the impact of noise on NeSpoF with a synthetic scene with varying noise levels. Figure 12(a) shows that NeSpoF results in a minor PSNR drop of 0.58 dB for the moderate noise level of Gaussian-Poisson noise with a standard deviation of 16. For the excessive noise level of standard deviation 80, the PSNR significantly drops by 7.07 dB in PSNR.
Material Dependency
The complexity of view-dependent spectro-polarimetric variation could vary with material types. In Figure 13, we evaluate NeSpoF on a synthetic scene consisting of multiple spheres with diverse materials. The experiment shows varying PSNRs depending on material types. Developing a material-robust representation for modeling the complex interplay between light and spectro-polarimetric BRDFs of real-world materials [Baek et al., 2020] remains a topic for future work.
Figure 13. Impact of material appearance on NeSpoF. We used 12 materials from the multispectral pBRDF dataset [Baek et al., 2020].
Image Resolution
NeSpoF can handle high-resolution input images. Figure 12 shows an increase of PSNR by 1.73 dB when using 1024$\times$1024 resolution compared to 512$\times$512 resolution. However, we note that capturing such high-resolution images with a low level of measurement noise could be challenging.
Number of Training Views
The performance of NeSpoF depends on the number of training views. Figure 12(b) shows that using 13 and 52 views for training results in the PSNRs of 28.82 dB and 34.98 dB, respectively. Improving NeSpoF for fewer views is an important problem, given the difficulty of capturing hyperspectral-polarimetric images at multiple views.